03 July 2014 UT Mini-queue night

Summary

Started the night with LBC in moderately clear sky and 0.8" seeing. Successfully observed 2 sets of V and r-sloan flats in LBCB and LBCR respectively. Started observations for Pal-5, but unfortunately thick clouds moved soon after twilight. Persevered through clouds and variable seeing for a bit until threatening conditions closed the dome. We remained closed for remainder of night. We consider the night lost to weather.

Plan

Start with LBC on high-priority UVa program (failed due to clouds). Switch to MODS for high-priority OSURC program (failed due to clouds). Turnover to Pathfinder around 12:30.

Log


02:50 Clearing straight above us. Opening.

03:09 Sky flats: V-Bessel on LBCB and r'-sloan on LBCR

All between ~20,000-10,000 counts. May have been light cirrus; should be checked carefully

03:21 Slewing to Pal-5

03:29 DOFPIA for Pal-5
Seeing estimate 0.86"on DIMM

03:36 Co-pointing for Pal-5

03:39 Starting Pal-5 (3) science OB ; unhappy with blue images (focus, elongation, etc) so we'll redo DOFPIA

03:50 DOFPIA again - now we have clouds and no pupils!

....waiting to see if clouds move off

04:00 Hole in the clouds. DOFPIA again

04:12 Started Pal-5 (3) again through clouds; Variable seeing between 1.0"-1.6".

Unfortunately everything after (B42209), (R42153) are strongly affected by clouds. Continuing in the hopes of a hole, but the clouds are
getting very thick now across the whole sky. A few images at the end may be good.

04:41 Running DOFPIA again; no stars

04:56 Still heavy clouds...

05:05 Slew to co-pointing field for NGC5194a (OSU Monitor)

05:10 DOFPIA -- some pupils; then none. Clouds again. trying to wait it out.

05:15 Tried a different field (Pal 3 again) with LBC, but aborted when we saw the All-Sky image. Images R52037 and B52035 are BAD.

05:20 Too many clouds for LBC collimation. Switching to MODS...

05:41 Checking pointing and trying collimation...very cloudy. Realizing now that it is safter to close up! Some rain reported soon after. Vent gates also closed.

06:00 Olga tries MODS test to see the behavior of AGW filter and central wavelength during Blue only observations centered ~4000. As expected,
the DD reports the correct filter (3) was inserted in the AGW automatically when the script was sent, but the central wavelength was not
updated. See IT 5177.

06:30 Olga tries same test but with fake guide star. (Restarted GCS to get fake guide star); same result. No new central wavelength.

06:45 Continued weather downtime...

07:15 Turning over to PATHFINDER closed dome.

Notes

  • The dome was closed due to clouds and there were few rain drops too (LBC run).

  • LBC handed over the telescope to Pathfinder by 07:27 hours.
  • Dome opened by 07:45 hours...
  • we are operating only the DX side of the telescope.
  • fast varying clouds. Can hardly see any star.
  • No PIE (on the ASM) as of now (07:59 Hours)
  • SE1, SE4, SE6 and SE10 with cheerios positioned in the GWS FoV to form a cross pattern:
    SE01 0.0mm, -72.0mm
    SE04 -72.0mm, 0.0mm
    SE06 0.0mm, +72.0mm
    SE10 +72.0mm, 0.0mm
    With the plate scale 0f 613 micrometer/arcsec a separation of 72mm corresponds to ~117.5 arcsec.

08:02 The telescope presetting to Vega.

08:05 Moved M3 and M2 to the night positions of Pathfinder settings.

08:08 The preset cancelled due to danger of rain.

08:13 Dome closed.


Waiting for clouds to vanish, better weather and stars in the sky to be visible...

08:42 Start of a couple of test presets with Dome closed

08:42 Telescope presetting to Dome Preset [test.catalogue].

08:47 Telescope presetting to Dome preset [Simple_2stars_2.catalogue].

08:50 Telescope Presetting to Dome preset [Simple_2stars_4.catalogue]. All these three presets are done with 'track' mode.

08:52 Telescope presetting to Dome Preset [Simple_2stars_4.catalogue]. Mode = 'track' and rotmode = 'PARALLACTIC'

08:55 The GWS bearing service is working. However, the rotmode has to be the default mode, 'Position Angle'

09:00 New dome preset sent with the rotmode = 'position angle'.

09:01 The bearing rotation service restarted. The bearing rotates compensating the sky rotation.

09:28 The sky looks better. Still we will wait for some more time. 09:30 Closed Dome Preset tests completed. Preset cancelled. Telescope set to Zenith and azimuth of 120.6 (nominal default position, allowing the telescope to drain the moisture if necessary)

09:42 Dome Opened. Clouds are clearing out. The weather is better.

09:50 Telescope presetting to Vega (mode = 'track', rotmode = 'Position Angle')

09:51 Bearing rotation started. Bearing start angle is 0 degrees.

09:53 M2 and M3 moved to the respective pathfinder nominal positions.

09:57 checking if the DSLR is getting the image of Vega. (the fiber is removed from its position earlier today, light should reach the DSLR without obstruction). We have some light, not sure if it is the stray light from dome.

10:03 taking different pictures in the DSLR changing the exposure and focus.

10:13 The star image is not yet seen at DSLR. There is a doubt if the tertiary mirror is working properly. So changing its position in steps and taking images to see if there are changes.

10:23 The dome lights were switched on for few mins to get light on the DSLR to confirm that the live view works. M3 was not in the right position. Therefore changed accordingly.

10:30 After adjustments, light on the DSLR.

10:35 The telescope is moved in elevation by 5 arcsec to check if the light is moved in the right direction. It does move, but may not be right amount. we therefore use Cheerios. We could not see the light on the cheerios. It might be due to the fact that DSLR and Telescope are out of focus.

10:52 We stop the bearing rotation and the change the preset so that the rotmode is 'Parallactic'

10:53 the telescope is moved in Det X / Det Y so that the star is on axis and properly focussed. - Applied pointing correction: (Det X, Det Y) = (-47,-62) - Mirror Positions: M3 Tip: -445 M3 Tilt: -200 Selector: 18.67 - Hexapod position ? 11:01 Offseting in Det X by 118 arcsec. Trying to see the star in the Cheerio.

11:04 Star image on the Cheerio of SE4. This is attained my moving the telescope and focussing.

11:09 Moving the star image from the Cheerio to inside the SE. Light on the SE and therefore light in the 4 pupils is visible. From the pupils, we could speculate that there is some astigmatism (the star image might not be completely inside the SE).

11:15 moving the telescope in det x / det y such that Vega is moved from on axis to 4 off axis positions such that the star image falls onto each of the 4 Chereios or SEs in the 72mm cross pattern.

(0, 0 ) DSLR (on-axis) (0, 117) SE1 (0,-117) SE6 The telescope is moved in Det Y and checking the position and brightness of the star image at the Cheerios. Iterating. (-117,0) SE10 ( 117,0) SE4 The star image offset pattern on the four Cheerios seem to be symmetrical, but slightly rotated wrt. the 72 mm pattern of the SEs.

11:43 The bearing is roatated by -2 degrees. Checking the 117 arcsec star image offset pattern again. The the pupil images appear on the 4 SEs (better than previous). Looks good, even though not illuminated completely.

11:47 preset again on Vega but with rotmode = 'Position Angle'. The bearing rotation service is activated. The bearing start angle is 0 deg.

11:51 Spiral search with a pitch of 7 arcsec to find Vega again.

12:05 Found Vega during spiral serch. Manual pointing correction to get Vega centered: offset (Ra,Dec) = (33,-119). Reason for the large pointing deviation unclear.

12:14 Telescope delta offsets towards SE4 (Ra,Dec) = (+94,+68) - now in position angle mode wrt the corrected pointing. Vega moves off the SLR in the correct direction. However, due to sunlight there is no chance to detect it anymore on the Cheerio nor on WFS CCD. positions are, (-117,+3), (+95,+68). These values are added to (33,-119), where the star is at the center of the DSLR.
Topic revision: r4 - 04 Jul 2014, ThomasBertram
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback